OpenStudy (anonymous):

Simplify using your knowledge of properties. 2(a + 4) + -2a + -8 Simplify using your knowledge of properties. 2(a + 4) + -2a + -8 @Mathematics

OpenStudy (anonymous):

0

OpenStudy (anonymous):

2(a + 4) + -2a + -8 = 2a+8-2a-8 =0

OpenStudy (anonymous):

2(a + 4) + -2a + 8 = 2a + 8 -2a -8 = 2a - 2a + 8 - 8 = 0